OpenStudy (anonymous):

The volume V(in liters) of 3 g of CO2 at 27C is given in terms of its pressure p(in atmospheres) by the formula V=1.68/p. What is the rateof change of V with respect to p when p=2(atm)? (Suggestion: use the fact that the derivative of f(x)=c/x is f'(x)= -c/x^2 if c is a constant; You can establish this by using the definition of the derivative.)

OpenStudy (anonymous):

v'(p) = dV/dp = -1.68 / p^2 v'(2) = -1.68 / 4 = 0.42